Kevin and Randy have a jar containing 33 coins, all of which are either quarters or nickels. The total value of the coins in the jar is $4.65. How many of each type of coin do they have?
Answer by richwmiller(17219)

x+y=33,
0.05*x+0.25*y=4.65
x=33-y
0.05*(33-y)+0.25*y=4.65
1.65-0.05y+0.25*y=4.65
0.2*y=3.0
y=15.0
x=33-y
x=18 at 0.05 and y=15 at 0.25
check
18+15=33
0.05*x+0.25*y=4.65
0.05*18+0.25*15=4.65
0.9+3.75=4.65
4.65=4.65
ok